OpenStudy (anonymous):

What are the real or imaginary solutions of the polynomial equation? 27x^3+125=0 I don't even know where to start and I have no idea how to do this

OpenStudy (amistre64):

subtract, divide ....

OpenStudy (anonymous):

what do you mean?

OpenStudy (amistre64):

i mean that if you are traveling down i60 and hit the mcdonalds in ohio .... what do you mean what do i mean?

OpenStudy (amistre64):

you can eithe rsolve for x, or recall a formula for the sum of cubes

OpenStudy (anonymous):

when I say what do you mean Im asking what d you divide and subtract

OpenStudy (anonymous):

\[27x ^{3}+125=0\] You want to solve for x so that when you plug this value in it will yield 0. Can you solve for x?

OpenStudy (anonymous):

subtract 125 and divide by 27?

OpenStudy (anonymous):

YES!

OpenStudy (amistre64):

that it :)

OpenStudy (anonymous):

Try it

OpenStudy (anonymous):

x^3=4.63

OpenStudy (anonymous):

then are you supposed to get the cube root of 4.63

OpenStudy (amistre64):

should be negative ... but then cbrt it for the real root

OpenStudy (anonymous):

so x=-1.67

OpenStudy (amistre64):

in general:\[(a^3+b^3)=(a+b)(a^2-ab+b^2)\]

OpenStudy (amistre64):

sp the real root is x = cbrt(-125/27) = -5/3

OpenStudy (anonymous):

\[27x ^{3}+125=0\] \[27x ^{3}=-125\] \[x ^{3}=-\frac{ 125 }{ 27 }\] \[x ^{3}=\sqrt[3]{-\frac{ 125 }{ 27 }}\]

OpenStudy (anonymous):

amistre what number is a and what is b

OpenStudy (anonymous):

which can be rewritten as \[x ^{3}=\frac{ \sqrt[3]{-125} }{ \sqrt[3]{27} }\]

OpenStudy (anonymous):

which number when you multiply it 3 times gives you -125??? Which number when you multiply 3 times gives you 27???

OpenStudy (amistre64):

now the roots for: 9x^2 - 15x +5 should be complex and assuming i placed the 3 and 5 them correctly

OpenStudy (amistre64):

a = cbrt(27x^3) and b= cbrt(125)

OpenStudy (anonymous):

sorry it's suppose to be \[x=\frac{ \sqrt[3]{-125} }{ \sqrt[3]{27} }\]

OpenStudy (anonymous):

so you get \[x=\frac{ -5 }{ 3 }\]

OpenStudy (anonymous):

sorry its taking me a while to reply openstudy is being really buggy for me

OpenStudy (anonymous):

If you plug this into the original equation \[27(-\frac{ 5 }{ 3 })^{3}+125=27(-\frac{ 125 }{ 27 })+125=-125+125=0\]

OpenStudy (anonymous):

so it checks out

OpenStudy (anonymous):

your solution is a real solution since you don't end up with any imaginary roots...that is we can take the cubic root of a negative number

OpenStudy (amistre64):

yeah, its been buggy for awhile now. comes and goes lately